### v1

@@ 5.1.2 Theoretical results
Alternatively, a fully supervised training strategy based on minimizing an empirical approximation to (5.33) gives similar convergence guarantees in theory [277, Thm. 4.1, p. 7], but empirically demonstrates worse performance than the unsupervised approach. However, suppose the Tikhonov regularizer is replaced by a more general regularizer, such as a sparsity-promoting one [278]. In that case, the supervised learning strategy becomes necessary because closed-form expressions for the optimal $(h^{\star},B^{\star})$ are no longer available. The resulting finite-sample analysis also becomes more difficult, requiring advanced tools from statistical learning theory [280].
@@
which delivers simultaneous diagonalizability with the normal operators corresponding to $\mathcal{A}$. The eigenvalues $\{\varsigma_{j}\}_{j\in\mathbb{N}}$ and $\{\delta^{2}\gamma_{j}\}_{j\in\mathbb{N}}$ from (5.36) are nonincreasing sequences. We assume that $\gamma_{1}=1$ [138, Sec. 2.2]. The parameter $\delta>0$ in (5.36) serves as a noise level. It will be convenient to define $\xi\coloneqq\eta/\delta$. The covariance operator of $\xi$ is $\sum_{j=1}^{\infty}\gamma_{j}\,e_{j}\otimes_{\mathcal{Y}}e_{j}$. We also assume that $\operatorname{\mathbb{E}}u=\overline{u}=0$.
@@
Under the simultaneous diagonalizability assumption (5.36) and the condition $\overline{u}=h^{\star}=0$, by comparing to [277, Eqn. (12), p. 5] when written in the SVD bases (5.35) of $\mathcal{A}$, we find that $\mathcal{R}_{\psi(\delta)}=\mathcal{R}_{h^{\star},B^{\star}}^{\Gamma}$ [138, Remark 2, p. 49]. That is, the optimal spectral regularizer—which did not enforce an explicit smoothness penalization—is identical to the optimal Tikhonov regularizer, which does enforce a quadratic penalty. It is interesting that Tikhonov structure emerges in the optimal spectral reconstruction as a consequence of the additive noise model (5.30) and the expected squared reconstruction error objective function (5.39). This finding is consistent with known results for finite-dimensional nonlinear regression [129, 281].
A heuristic explanation for this phenomenon in the current linear infinite-dimensional setup is as follows. The optimal generalized Tikhonov reconstruction (5.32) implicitly computes the MAP point of a Gaussian posterior $\mu^{y}$ determined by a Gaussian likelihood $\mathcal{N}(\mathcal{A}u,\Gamma)$ and Gaussian prior $\mathcal{N}(0,\Sigma)$, which match the first two moments of $\operatorname{Law}(y\,|\,u)$ and $\operatorname{Law}(u)$, respectively. This immediately shows that the optimal Tikhonov reconstruction is the posterior mean $\operatorname{\mathbb{E}}_{u\sim\mu^{y}}[u]$ of $\mu^{y}$, agreeing with known formulas [67, Prop. 3.1, p. 2630]. To relate back to the optimal spectral regularizer, one would have to argue that since $\mathcal{R}_{\psi}$ is linear in $\psi$ and $y$ is linearly related to $u$ in expectation, the minimizer of (5.39) over $\psi$ (i.e., linear maps) should equal the Gaussian posterior mean instead of the Bayes estimator $\operatorname{\mathbb{E}}[u\,|\,y]$. However, a proper proof based on this intuition in the present infinite-dimensional setting is beyond the scope of the chapter.
We close this subsection by summarizing accuracy guarantees for the optimal spectral regularizer $\mathcal{R}_{\psi(\delta)}$. The preceding discussion implies that consistency of learned spectral regularization is equivalent to consistency of the Gaussian posterior mean estimator. Consistency of the latter is well known [67, 69]. Under mild distributional assumptions, more direct arguments [279, Thm. 1, p. 1351] establish the pointwise average-case convergence
@@
Compare this result with (4.36), which gives pointwise worst-case convergence. The convergence in (5.41) is actually uniform over certain bounded sets of noise distributions with noise level at most $\delta$ and covariance $\Gamma$; see [138, Eqn. (2.8), p. 48]. Going beyond pointwise error, we also have the full mean square convergence
@@
By using well-established Bayesian theory, it should also be possible to obtain quantitative convergence rates as $\delta\to 0$ under suitable decay conditions on the singular values of all the involved operators [67]. The misspecified case, in which the noise distribution in the training data possibly differs from the one used to compute the squared reconstruction error, has also been handled [138]. In all cases, the diagonal analysis suggests that choosing the training data to be corrupted by white noise (i.e., $\gamma_{j}=1$ for all $j$) leads to better convergence properties for solving the inverse problem. This message is consistent with a related insight for forward linear operator learning in the following sense. In [148, Corollary 3.6, pp. 17–18], it was observed that training data $\{y_{n}\}$ corrupted with white noise worsens convergence rates for learning the forward operator. Since the forward problem becomes harder, the inverse problem intuitively becomes easier as claimed.
Beyond the mean square optimal estimator (5.40), the preceding spectral regularization framework has also been combined with plug-and-play estimators [138, Sec. 4]. We will overview such plug-and-play methods next.
@@ 5.2 Denoising networks and plug-and-play methods
A major breakthrough in bridging model-based and learning-based approaches to inverse problems was the realization that powerful denoisers—including neural networks trained to denoise corrupted signals—implicitly encode high-quality priors. Plug-and-play (PnP) methods capitalize on this insight by integrating pre-trained image denoising networks into iterative optimization algorithms as implicit regularizers. Instead of explicitly specifying a prior through a regularization functional $R$, PnP methods replace the proximal step in classical algorithms such as the alternating direction method of multipliers (ADMM) or proximal gradient descent with a learned denoiser $\mathcal{D}$. In the following, we briefly review the latest developments in PnP. Section 5.2.1 introduces the PnP framework and Section 5.2.2 surveys existing theoretical guarantees for the framework.
@@ 5.2.2 Theoretical analysis of PnP
A complementary line of work reframes PnP as a probabilistic inference method. Bouman et al. [289] introduce Generative PnP (GPnP), which transforms deterministic PnP iterations into a Markov chain by replacing each update with a stochastic proximal generator. The resulting dynamics sample from the Bayesian posterior distribution, thus providing both point estimates and uncertainty quantification. Wu et al. [38] demonstrate that any inverse problem can be reformulated as a Gaussian image denoising task. This allows one to embed pretrained diffusion models within lightweight MCMC schemes, thereby integrating score-based generative models into the PnP framework. In this setting, denoisers serve as approximations to the score function. They enable posterior sampling via Langevin dynamics or stochastic proximal steps.
@@ 6 Conclusion and outlook
*Acknowledgments. The authors thank Maarten de Hoop, Bamdad Hosseini, Matti Lassas, Youssef Marzouk, Richard Nickl, and Andrew Stuart for numerous discussions that helped to shape this chapter. The authors are also grateful to Nicolas Guerra for helpful comments on a previous version of the manuscript.

### v2

@@ 5.1.2 Theoretical results
Alternatively, a fully supervised training strategy based on minimizing an empirical approximation to (5.36) gives similar convergence guarantees in theory [315, Thm. 4.1, p. 7], but empirically demonstrates worse performance than the unsupervised approach. However, suppose the Tikhonov regularizer is replaced by a more general regularizer, such as a sparsity-promoting one [316]. In that case, the supervised learning strategy becomes necessary because closed-form expressions for the optimal $(h^{\star},B^{\star})$ are no longer available. The resulting finite-sample analysis also becomes more difficult, requiring advanced tools from statistical learning theory [318].
@@
which delivers simultaneous diagonalizability with the normal operators corresponding to $\mathcal{A}$. The eigenvalues $\{\varsigma_{j}\}_{j\in\mathbb{N}}$ and $\{\delta^{2}\gamma_{j}\}_{j\in\mathbb{N}}$ from (5.39) are nonincreasing sequences. We assume that $\gamma_{1}=1$ [148, Sec. 2.2]. The parameter $\delta>0$ in (5.39) serves as a noise level. It will be convenient to define $\xi\coloneqq\eta/\delta$. The covariance operator of $\xi$ is $\sum_{j=1}^{\infty}\gamma_{j}\,e_{j}\otimes_{\mathcal{Y}}e_{j}$. We also assume that $\operatorname{\mathbb{E}}u=\overline{u}=0$.
@@
Under the simultaneous diagonalizability assumption (5.39) and the condition $\overline{u}=h^{\star}=0$, by comparing to [315, Eqn. (12), p. 5] when written in the SVD bases (5.38) of $\mathcal{A}$, we find that $\mathcal{R}_{\psi(\delta)}=\mathcal{R}_{h^{\star},B^{\star}}^{\Gamma}$ [148, Remark 2, p. 49]. That is, the optimal spectral regularizer—which did not enforce an explicit smoothness penalization—is identical to the optimal Tikhonov regularizer, which does enforce a quadratic penalty. It is interesting that Tikhonov structure emerges in the optimal spectral reconstruction as a consequence of the additive noise model (5.33) and the expected squared reconstruction error objective function (5.42). This finding is consistent with known results for finite-dimensional nonlinear regression [139, 319].
A heuristic explanation for this phenomenon in the current linear infinite-dimensional setup is as follows. The optimal generalized Tikhonov reconstruction (5.35) implicitly computes the MAP point of a Gaussian posterior $\mu^{y}$ determined by a Gaussian likelihood $\mathcal{N}(\mathcal{A}u,\Gamma)$ and Gaussian prior $\mathcal{N}(0,\Sigma)$, which match the first two moments of $\operatorname{Law}(y\,|\,u)$ and $\operatorname{Law}(u)$, respectively. This immediately shows that the optimal Tikhonov reconstruction is the posterior mean $\operatorname{\mathbb{E}}_{u\sim\mu^{y}}[u]$ of $\mu^{y}$, agreeing with known formulas [71, Prop. 3.1, p. 2630]. To relate back to the optimal spectral regularizer, one would have to argue that since $\mathcal{R}_{\psi}$ is linear in $\psi$ and $y$ is linearly related to $u$ in expectation, the minimizer of (5.42) over $\psi$ (i.e., linear maps) should equal the Gaussian posterior mean instead of the Bayes estimator $\operatorname{\mathbb{E}}[u\,|\,y]$. However, a proper proof based on this intuition in the present infinite-dimensional setting is beyond the scope of the chapter.
We close this subsection by summarizing accuracy guarantees for the optimal spectral regularizer $\mathcal{R}_{\psi(\delta)}$. The preceding discussion implies that consistency of learned spectral regularization is equivalent to consistency of the Gaussian posterior mean estimator. Consistency of the latter is well known [71, 73]. Under mild distributional assumptions, more direct arguments [317, Thm. 1, p. 1351] establish the pointwise average-case convergence
@@
Compare this result with (4.36), which gives pointwise worst-case convergence. The convergence in (5.44) is actually uniform over certain bounded sets of noise distributions with noise level at most $\delta$ and covariance $\Gamma$; see [148, Eqn. (2.8), p. 48]. Going beyond pointwise error, we also have the full mean square convergence
@@
By using well-established Bayesian theory, it should also be possible to obtain quantitative convergence rates as $\delta\to 0$ under suitable decay conditions on the singular values of all the involved operators [71]. The misspecified case, in which the noise distribution in the training data possibly differs from the one used to compute the squared reconstruction error, has also been handled [148]. In all cases, the diagonal analysis suggests that choosing the training data to be corrupted by white noise (i.e., $\gamma_{j}=1$ for all $j$) leads to better convergence properties for solving the inverse problem. This message is consistent with a related insight for forward linear operator learning in the following sense. In [158, Corollary 3.6, pp. 17–18], it was observed that training data $\{y_{n}\}$ corrupted with white noise worsens convergence rates for learning the forward operator. Since the forward problem becomes harder, the inverse problem intuitively becomes easier as claimed.
Beyond the mean square optimal estimator (5.43), the preceding spectral regularization framework has also been combined with plug-and-play estimators [148, Sec. 4]. We will overview such plug-and-play methods next.
@@ 5.2 Denoising networks and plug-and-play methods
A major breakthrough in bridging model-based and learning-based approaches to inverse problems was the realization that powerful denoisers—including neural networks trained to denoise corrupted signals—implicitly encode high-quality priors [320]. Plug-and-play (PnP) methods capitalize on this insight by integrating pre-trained image denoising networks into iterative optimization algorithms as implicit regularizers. Instead of explicitly specifying a prior through a regularization functional $R$, PnP methods replace the proximal step in classical algorithms such as the alternating direction method of multipliers (ADMM) or proximal gradient descent with a learned denoiser $\mathcal{D}$. In the following, we briefly review the latest developments in PnP. Section 5.2.1 introduces the PnP framework and Section 5.2.2 surveys existing theoretical guarantees for the framework.
@@ 5.2.2 Theoretical analysis of PnP
A complementary line of work reframes PnP as a probabilistic inference method. Bouman et al. [328] introduce Generative PnP (GPnP), which transforms deterministic PnP iterations into a Markov chain by replacing each update with a stochastic proximal generator. The resulting dynamics sample from the Bayesian posterior distribution corresponding to the prior specified implicitly by the denoiser, and thus can provide both point estimates and uncertainty quantification. Wu et al. [40] demonstrate that any inverse problem can be reformulated as a Gaussian image denoising task. This allows one to embed pretrained diffusion models within lightweight MCMC schemes, thereby integrating score-based generative models into the PnP framework. In this setting, denoisers serve as approximations to the score function. They enable posterior sampling via Langevin dynamics or stochastic proximal steps.
@@ 6 Conclusion and outlook
*Acknowledgments. The authors thank Maarten de Hoop, Bamdad Hosseini, Matti Lassas, Youssef Marzouk, Richard Nickl, and Andrew Stuart for numerous discussions that helped to shape this chapter. The authors are also grateful to Nicolas Guerra and two anonymous referees for helpful comments on a previous version of the manuscript.